OpenStudy (anonymous):

integral of (3x-2)/(x+1) using partial fractions

OpenStudy (anonymous):

you can divide the numerator by denominator or you can do like this. \[\frac{ 3x-2 }{x+1 }=3+\frac{ 3\left( -1 \right)-2 }{x+1 }=3-5\frac{ 1 }{ x+1 }\] \[\int\limits \frac{ 3x-2 }{x+1 }dx=\int\limits3 dx-5\int\limits \frac{ 1 }{x+1 }dx+c=3x-5\ln \left( x+1 \right)+c\]